# Solving nonconvex Hamilton–Jacobi–Isaacs equations with PINN-based policy iteration

### v1

@@ 4.1 Implementation Setup
To quantitatively evaluate the accuracy of the learned value functions, we compute reference solutions to the HJI equation using an explicit finite difference method on a uniform grid. We employ backward time integration with central differences in space, and discretize the diffusion term using a second-order scheme, with homogeneous Neumann boundary conditions imposed on the spatial domain. To maintain stability in the presence of diffusion, the time and space steps are chosen such that $\Delta t=\Delta x^{2}$. The resulting numerical solutions approximate the viscosity solution and are used to compute the relative $L^{2}$ errors.
@@ 4.2 Two-dimensional optimal path planning with a moving obstacle
$\begin{cases}\mathrm{d}X(s)&=(a(s)+b(s))\mathrm{d}t+\sigma\mathrm{d}W_{t},\quad\text{for}\quad s\in(t,T],\\ X(t)&=x\in\mathbb{R}^{2}\end{cases}$
where $a(s)$ satisfying $|a(s)|\leq 1$ denotes the control input of the robot (Player I), $b(s)$ satisfying $|b(s)|\leq\delta$ for some $\delta>$ represents the adversarial disturbance (Player II), and $\sigma\in\mathbb{R}^{2\times 2}$ is the noise matrix. Given $x_{\text{goal}}\in\mathbb{R}^{2}$ and weights $\lambda_{i}$ for $i=1,2,3$, the cost functional is defined as:
$J(t,\bm{x};a,b)=\mathbb{E}[\int_{0}^{T}(\lambda_{1}|a(s)|^{2}+\lambda_{2}\phi(s,X(s)))\mathrm{d}s+\lambda_{3}|X(T)-x_{\mathrm{goal}}|^{2}],$
@@
$\phi(s,x)=\exp(-\frac{\|x-x_{\mathrm{obs}}(s)\|^{2}}{2\varepsilon^{2}}),\quad x_{\mathrm{obs}}(s)=\begin{bmatrix}0.5\cos(\pi s)\\ 0.5\sin(\pi s)\end{bmatrix}.$
@@
$v(t,x)=\sup_{\beta\in\Gamma_{t}}\inf_{a\in\mathcal{A}_{t}}\mathbb{E}[\int_{t}^{T}c(s,x(s),a(s),\beta[a](s))\mathrm{d}s+g(x(T))],$
@@
$H(t,x,p)=\sup_{\textbf{b}\in B}\inf_{\textbf{a}\in A}[\lambda_{1}|\textbf{a}|^{2}+\lambda_{2}\phi(t,x)+p(\textbf{a}+\textbf{b})].$
@@ 5 Conclusion
Furthermore, unlike direct one-shot approaches, our iterative framework enables systematic quantification of approximation errors. In particular, the $L^{2}$ error between the learned and true value functions can be rigorously bounded even in nonconvex settings, owing to the linearized structure of each policy evaluation step. Nevertheless, the method inherits limitations intrinsic to PINN-based approaches, such as sensitivity to network initialization and challenges in gradient stability over long training horizons.
An important limitation of the present work is the assumption of non-degenerate diffusion. In the absence of noise, the problem becomes a first-order Hamilton–Jacobi–Isaacs equation arising in deterministic differential games. This setting lacks the regularizing properties of second-order terms and poses distinct theoretical and computational challenges. We leave the development of PINN-based policy iteration schemes for such first-order HJI problems to future work.
@@ Appendix B Proof of Value Function Decomposition
Let $x=[x_{0},x_{1},\dots,x_{N-1}]^{\top}\in\mathbb{R}^{N}$ be the full state vector, where $x_{0}$ is the publisher state and $x_{i}$ is the $i$-th subscriber state. The terminal cost is given by a sum of pairwise costs:
@@
where $P_{i}:\mathbb{R}^{N}\to\mathbb{R}^{2}$ denotes the projection $P_{i}x=[x_{0},x_{i}]^{\top}$.

### v2

@@ 4.1 Implementation Setup
To quantitatively evaluate the accuracy of the learned value functions, we compute reference solutions to the HJI equation using an explicit finite difference method on a uniform grid. We employ backward time integration with central differences in space, and discretize the diffusion term using a second-order scheme, with homogeneous Neumann boundary conditions imposed on the spatial domain. To maintain stability in the presence of diffusion, the time and space steps are chosen such that $\Delta t=\Delta x^{2}$. The resulting numerical solutions approximate the viscosity solution and are used to compute the relative $L^{2}$-errors.
@@ 4.2 Two-dimensional optimal path planning with a moving obstacle
$\begin{cases}\mathrm{d}X(s)&=(a(s)+b(s))\mathrm{d}s+\sigma\mathrm{d}W_{s},\quad\text{for}\quad s\in(t,T],\\ X(t)&=x\in\mathbb{R}^{2}\end{cases}$
where $a(s)$ satisfying $|a(s)|\leq 1$ denotes the control input of the robot (Player I), $b(s)$ satisfying $|b(s)|\leq\delta$ for some $\delta>0$ represents the adversarial disturbance (Player II), and $\sigma\in\mathbb{R}^{2\times 2}$ is the noise matrix. Given $x_{\text{goal}}\in\mathbb{R}^{2}$ and weights $\lambda_{i}$ for $i=1,2,3$, the cost functional is defined as:
$J(t,x;a,b)=\mathbb{E}\left[\int_{0}^{T}(\lambda_{1}|a(s)|^{2}+\lambda_{2}\phi(s,X(s)))\mathrm{d}s+\lambda_{3}|X(T)-x_{\mathrm{goal}}|^{2}\right],$
@@
$\phi(s,x)=\exp\left(-\frac{\|x-x_{\mathrm{obs}}(s)\|^{2}}{2\varepsilon^{2}}\right),\quad x_{\mathrm{obs}}(s)=\begin{bmatrix}0.5\cos(\pi s)\\ 0.5\sin(\pi s)\end{bmatrix}.$
@@
$v(t,x)=\sup_{\beta\in\Gamma_{t}}\inf_{a\in\mathcal{A}_{t}}\mathbb{E}\left[\int_{t}^{T}c(s,x(s),a(s),\beta[a](s))\mathrm{d}s+g(x(T))\right],$
@@
$H(t,x,p)=\sup_{\textbf{b}\in B}\inf_{\textbf{a}\in A}[\lambda_{1}|\textbf{a}|^{2}+\lambda_{2}\phi(t,x)+p\cdot(\textbf{a}+\textbf{b})].$
@@ 5 Conclusion
Furthermore, unlike direct one-shot approaches, our iterative framework enables systematic quantification of approximation errors. In particular, the $L^{2}$-error between the learned and true value functions can be rigorously bounded even in nonconvex settings, owing to the linearized structure of each policy evaluation step. Nevertheless, the method inherits limitations intrinsic to PINN-based approaches, such as sensitivity to network initialization and challenges in gradient stability over long training horizons.
An important limitation of the present work is the assumption of non-degenerate diffusion. In the absence of noise, the problem becomes a first-order Hamilton–Jacobi–Isaacs equation arising in deterministic differential games. This setting lacks the regularizing properties of second-order terms and poses distinct theoretical and computational challenges. We leave the development of PINN-based policy iteration schemes for such first-order HJI problems to future work. Nonconvex HJ equations with degeneracy have seen significant progress, notably through policy iteration [9] and the nonlinear adjoint approach [20]. We expect that combining these ideas with our method could lead to effective extensions to degenerate or first-order problems.
@@ Appendix B Proof of Value Function Decomposition
Let $x=(x_{0},x_{1},\dots,x_{N-1})\in\mathbb{R}^{N}$ be the full state vector, where $x_{0}$ is the publisher state and $x_{i}$ is the $i$-th subscriber state. The terminal cost is given by a sum of pairwise costs:
@@
where $P_{i}:\mathbb{R}^{N}\to\mathbb{R}^{2}$ denotes the projection $P_{i}x=(x_{0},x_{i}).$